Two forces act on a 2-kg box. The first force is 10 N to the right, while the second force if 5 N to left. Determine the resultant acceleration. [Hint: Solve for the value of the net force applied to the box. Then, solve for the box’s acceleration. Remember to specify the direction of the acceleration when giving your final answer.]


1
Expert's answer
2021-11-02T10:00:07-0400

Let the direction to the right be positive. Let's first find the net force applied to the box:


"F_{net}=F_1-F_2=10\ N-5\ N=5\ N."

The sign plus means that the net force directed to the right.

Finally, we can find the box’s acceleration from the Newton's Second Law of Motion:


"F_{net}=ma,""a=\dfrac{F_{net}}{m}=\dfrac{5\ N}{2\ kg}=2.5\ \dfrac{m}{s^2}."

The sign plus means that the box’s acceleration directed to the right.